I want to achieve the 10mcg concentration per 60ml of juice

The powder spec is

2500mcg/g concentration
Molecular mass 384.65g/mol
1
Expert's answer
2020-12-14T14:44:56-0500

Target solution

mass = 10 microgram

Volume = 60ml

.: Concentration = 10/60 = 0.167microgram/gram


Powder stock

Concentration = 2500 microgram/gram


Using, "C_1V_1 = C_2V_2"

0.167 × 60 = 2500 × V2

V2 = 0.167×60/2500 = 0.004


This means that you take 0.004ml of the stock and put it in 59.996ml of water for your target solution to have a concentration of 0.167mcg/ml.